OpenStudy (anonymous):

can 2(1-sin(theta)) be the same as 2-sin 2(theta)

OpenStudy (dape):

Nope, the 2 must be "on" the sin, not inside it. If you put it inside you change the frequency of the sin-wave.

OpenStudy (dape):

For example, the period of \(\sin(2\theta)\) is \(\pi\), while the period of the usual \(\sin\theta\) is \(2\pi\).

OpenStudy (anonymous):

so when u distribute the 2 is it going to be 2-2sin(theta)

OpenStudy (dape):

Yes, that would be right!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how do you integrate \[\cos(\theta)\sin(\theta)\]